# Aligned Novel View Image and Geometry Synthesis via Cross-modal Attention Instillation

### v2

@@ A.1 Training details
In our training procedure, we initialize the image denoising U-Net from Stable Diffusion 2.1 and fine-tune on RealEstate10K [34], Co3D [22], and MVImgNet [33]. Reference networks share identical architecture and initial weights with the denoising U-Net but exclude timestep embeddings, serving solely for semantic feature extraction. Our model is trained using a batch size of 1 with gradient accumulation, employing mixed-precision training (bf16) to optimize memory usage and computational efficiency. We utilize 8-bit Adam optimizer with standard hyperparameters: $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, weight decay of 1×$10^{-2}$, and $\epsilon$ = 1×$10^{-8}$. The learning rate is set to 1×10 with a constant scheduler and minimal warmup of 1 step.
For multi-view training, we sample 4 viewpoints total with 3 reference views and 1 target view (index 1), maintaining a temporal margin of 30 frames between reference and target images to ensure sufficient viewpoint diversity. All training images are resized to 512×512 resolution. We employ the xFormers memory-efficient attention mechanisms to handle the computational demands of multi-view aggregated attention. The noise scheduler follows a scaled linear beta schedule with 1000 training timesteps, $\beta_{start}$ = 0.00085, $\beta_{end}$ = 0.012, and steps offset of 1, without sample clipping. Validation is performed every 2000 training steps to monitor convergence and prevent overfitting.
A.2 Implementation details
Geometry for correspondence conditioning is generated using VGGT [28], with only reference view pointmaps used for warping and proximity-based mesh conditioning. Ground truth geometry for MSE loss supervision is also predicted by VGGT using target images, enabling the model to maintain scale-and-shift alignment with reference geometry automatically. This approach allows exact reconstruction of known geometry while completing unknown regions in alignment with reference observations.
To accelerate and stabilize training, we separately fine-tune the pointmap denoising U-Net and its reference network, initializing from Marigold [13]’s normal prediction weights with identical multiview aggregated attention as the image U-Net. This separate initialization enables both branches to learn robust independent representations before cross-modal attention distillation, where geometry networks benefit from deterministic image cues, significantly improving prediction consistency. However, as discussed in our main paper, geometry-only training cannot achieve perfectly aligned predictions, necessitating cross-modal attention distillation for optimal performance.
@@ Appendix B Additional results
B.1 Additional scene completion results
@@
Figure 10: Qualitative results on single-view extrapolative setting. Our qualitative comparison of previous approaches demonstrates our model’s extrapolative capabilities to plausibly generate locations not seen in reference images while reconstructing faithfully the known regions.
@@
Figure 11: Qualitative comparison on two-view extrapolative setting. Our qualitative comparison of previous approaches demonstrates our model’s extrapolative capabilities to plausibly generate locations not seen in reference images while reconstructing faithfully the known regions.
@@
For two-view novel view synthesis, as demonstrated in Fig. 11, we conduct comprehensive evaluations across Co3D [22] datasets, demonstrating consistent performance improvements when additional reference information becomes available. The model effectively aggregates information from both reference views through our multi-view attention mechanism, resulting in more accurate geometry estimation and higher-fidelity image synthesis. Our architecture’s ability to seamlessly handle two-view inputs during inference, despite being trained on this configuration, validates the effectiveness of our correspondence conditioning and attention aggregation strategies.
B.2 Qualitative comparison with other generative models
Figure 12: Qualitative comparison with other warping-and-inpainting models. Our qualitative comparison of previous approaches demonstrates our model’s extrapolative capabilities to plausibly generate locations not seen in reference images while reconstructing faithfully the known regions.
Warping-and-inpainting models.
In Fig. 12, conduct qualitative evaluation against warping-and-inpainting methods using single reference images (LucidDreamer [5], GenWarp [25]) on the DTU dataset [11], which was excluded from training for all methods, demonstrating zero-shot generalization capabilities. For a fair comparison, all warping-and-inpainting methods utilize VGGT [28] as the shared geometry prediction model. Since our model excels in extrapolative scenarios, we focus our qualitative analysis on challenging extrapolative view selections that sample the most distant target cameras in the dataset, where substantial viewpoint changes test the limits of each method’s synthesis capabilities. Visual results in Fig. 12 demonstrate our model’s superior ability to effectively filter artifacts and noise from predicted point clouds during the warping process, producing clean, geometrically consistent images with well-preserved structural details. In stark contrast, baseline warping-and-inpainting methods produce numerous visual artifacts, blurred regions, and feature inconsistencies, particularly in occluded areas requiring substantial inpainting. These qualitative comparisons clearly highlight our approach’s robust handling of noisy geometric conditions and its capacity to maintain both photometric quality and geometric coherence across challenging viewpoint transformations.
B.3 Qualitative ablation
Figure 13: Ablation results. Our qualitative comparison of previous approaches demonstrates our model’s extrapolative capabilities to plausibly generate locations not seen in reference images while reconstructing faithfully the known regions.
B.4 Depth prediction accuracy
Method Image Geometry Geometry (Recon) Geometry (Inpainting)
[…]
Baseline 14.51 0.551 0.394 0.250 0.685 0.152 0.819 0.308 0.531
Ours 19.43 0.654 0.201 0.140 0.787 0.094 0.912 0.198 0.684
Table 5: Depth prediction accuracy. Incomplete, under experiment being filled.
We evaluate the depth prediction accuracy of our method using standard metrics such as PSNR, SSIM, and LPIPS. Quantitative comparisons against baseline methods demonstrate that our cross-modal attention instillation framework significantly reduces depth prediction error, particularly in challenging extrapolative scenarios. The integration of geometric cues via proximity-based mesh conditioning plays a critical role in refining the depth predictions, as evidenced by our ablation studies.
In addition to the numerical evaluations, our qualitative assessments reveal that the depth maps generated by our approach exhibit smooth transitions and accurate handling of occluded regions. The enhanced structural consistency is directly attributable to the attention instillation mechanism, which transfers reliable spatial attention maps from the image domain to the geometry networks. These results underline the effectiveness of our method in achieving high-quality depth estimation that closely aligns with the ground-truth geometry.

### v3

@@ A.1 Training details
[…]
For multi-view training, we sample 4 viewpoints total with 3 reference views and 1 target view (index 1), maintaining a temporal margin of 30 frames between reference and target images to ensure sufficient viewpoint diversity. All training images are resized to 512×512 resolution. We employ the xFormers memory-efficient attention mechanisms to handle the computational demands of multi-view aggregated attention. The noise scheduler follows a scaled linear beta schedule with 1000 training timesteps, $\beta_{start}$ = 0.00085, $\beta_{end}$ = 0.012, and a step offset of 1, without sample clipping. Validation is performed every 2000 training steps to monitor convergence and prevent overfitting.
A.2 Computational Efficiency
Despite employing a dual-branch architecture with multi-view attention mechanisms, our model maintains competitive inference speed and memory efficiency. For joint 2-view image and geometry generation, our method requires 9.81 seconds on an A6000 GPU with 28GB memory consumption. Notably, cross-modal attention instillation contributes to reduced memory overhead: the geometry branch reuses pre-computed attention maps from the image branch rather than computing separate attention operations, thereby saving both memory and computation. Furthermore, our image branch can operate independently for image-only novel view synthesis, requiring only 14GB memory and 4.3 seconds inference time—comparable to standard diffusion models. This flexibility allows users to trade off between full geometric consistency (dual-branch mode) and faster image-only generation depending on application requirements.
A.3 Implementation details
Geometry for correspondence conditioning is generated using VGGT Wang et al. (2025), with only reference view pointmaps used for warping and proximity-based mesh conditioning. Ground truth geometry for MSE loss supervision is also predicted by VGGT using target images, enabling the model to maintain scale-and-shift alignment with reference geometry automatically. This approach allows exact reconstruction of known geometry while completing unknown regions in alignment with reference observations.
To accelerate and stabilize training, we separately fine-tune the pointmap denoising U-Net and its reference network, initializing from Marigold Ke et al. (2024)’s normal prediction weights with identical multiview aggregated attention as the image U-Net. This separate initialization enables both branches to learn robust independent representations before cross-modal attention distillation, where geometry networks benefit from deterministic image cues, significantly improving prediction consistency. However, as discussed in our main paper, geometry-only training cannot achieve perfectly aligned predictions, necessitating cross-modal attention distillation for optimal performance.
@@ Appendix B Additional results
B.1 Additional Comparison Against Large Models
Figure 9: Qualitative comparison with large models on Navi (Jampani et al., 2023) Dataset. Our model shows competitive performance against large model-based novel view synthesis models.
We additionally compare against VGGT (Wang et al., 2025), LVSM (Jin et al., 2024) and ViewCrafter (Yu et al., 2024), large model-based novel view synthesis models. As shown in Fig. 9, our method excels in extrapolative scenarios with significant unobserved regions. LVSM fails beyond observable boundaries, producing blurry occluded content, while ViewCrafter generates geometrically degraded imagery under extrapolative settings and requires 209.19 seconds for 25 frames on an A6000 GPU. Our approach synthesizes plausible content in only 9.67 seconds on average, maintaining high-fidelity detail and geometric consistency, demonstrating superior quality and efficiency.
[…]
ViewCrafter 1-view 14.04 0.390 0.332
Ours (Single-view) 1-view 15.56 0.609 0.184
LVSM 2-view 15.23 0.499 0.415
Ours (Stereo-view) 2-view 15.58 0.615 0.184
Table 7: Comparison to large model baselines. We quantitatively compare our model against recent large-scale models.
Quantitative evaluation. To further validate our model’s generalization capability, we conduct a quantitative evaluation on the DTU dataset under extrapolative view settings in Table 7. We compare against ViewCrafter and LVSM under both single-view and two-view settings. Results demonstrate that our model achieves superior performance across all metrics in both configurations: under single-view input, our model outperforms ViewCrafter by +1.52 dB PSNR, +0.219 SSIM, and 0.148 LPIPS; under two-view input, our model surpasses LVSM by +0.32 dB PSNR, +0.116 SSIM, and 0.231 LPIPS. These quantitative results on DTU corroborate our qualitative findings on RealEstate10k and Navi datasets, confirming our model’s competitive performance against state-of-the-art methods across diverse benchmarks.
B.2 Robustness to perturbations in correspondence condition
Figure 10: Robustness experiments regarding geometric noise. Visualization of MoAI’s robustness against noise added to the predicted geometry.
In Fig. 10, we present qualitative experiments examining robustness to geometric noise in the predicted geometry. We apply varying levels of Gaussian noise (standard deviations of 6%, 10%, and 15% relative to the standard deviation of reference point cloud coordinates) to the point locations in the predicted pointmaps and use these noisy pointmaps as correspondence conditions. Despite high noise levels, our model exhibits minimal performance degradation, demonstrating strong robustness against errors and artifacts in the predicted geometry.
Figure 11: Robustness experiments regarding sparsity. Visualization of MoAI’s robustness against increased sparsity of predicted geometry.
In Fig. 11, we provide similar experiments regarding the sparsity of the predicted geometry: we randomly mask out a certain percentage of points (30%, 50% and 80%) from the predicted pointmap, and use this sparse pointmap as the correspondence condition. Despite high sparsity in the correspondence condition, our model shows minimal performance degradation even at 80% masking scenario, demonstrating robustness against sparsity in predicted geometry. These results confirm that our framework’s generative approach effectively mitigates errors from external priors, maintaining our pose-free methodology without compromising output quality.
B.3 Qualitative evaluation with in-the-wild data
We provide additional results demonstrating our model’s strong generalization to unseen domains, including in-the-wild and urban data. We conducted experiments on two new datasets from different domains: the CityScapes (Cordts et al., 2016) dataset, containing urban street-view multi-view images; the MegaDepth (Li and Snavely, 2018) dataset. None of these datasets was used during training, thereby validating our model’s generalization capabilities. Fig. 12 presents qualitative results demonstrating high-quality novel view synthesis and geometry reconstruction on in-the-wild data, confirming strong generative performance across diverse domains.
Figure 12: Qualitative results on in-the-wild / urban data. We provide generalization results on unseen in-the-wild / urban datasets (MegaDepth, CityScapes), for which our model is capable of conducting high-fidelity novel view synthesis.
B.4 Evaluation regarding geometric consistency in occlusion and shading
Figure 13: Occlusion handling in extrapolative viewpoints. We demonstrate the performance of our model under occlusion / lighting / shadow handling in extrapolative (greater than 90 degrees) target camera pose.
Fig. 13, in addition to Fig. 4, demonstrates extrapolative view cases where our model generates target views more than 90° away from the reference viewpoints, requiring generation of completely unseen regions. Our model maintains physical consistency when extrapolating beyond 90° view differences, including correct shading and occlusion handling.
B.5 Alignment evaluation with different external pose prediction model
Figure 14: Alignment visualization with different external model. We demonstrate alignment visualization between reference view geometry and generated target view geometry, evaluated with DepthAnything V3 (Lin et al., 2025).
The alignment of our generated geometry to the predicted geometry is inherently ensured by our architectural design, regardless of where the target camera pose derives from. Specifically, we formulate geometry generation as inpainting (or completion) of the pointmap that has already been projected to the target camera pose. In this formulation, the model learns to leave the projected geometry unchanged and simply generate extended geometry around it (similar to the image inpainting task), thereby maintaining alignment within the projected geometry according to the given camera pose. Therefore, geometry completion and alignment take place agnostic to the target-camera normalization within our architecture.
To demonstrate this, we provide a cross-predictor validation experiment in Fig. 14, where we employ an alternative external geometry model (DepthAnything v3 (Lin et al., 2025)) at evaluation time to assess alignment between the externally predicted geometry and our model’s generated geometry. The visualized results show that our model achieves accurate alignment and high-quality geometry completion regardless of which external prediction model is used for geometry and pose prediction.
Figure 15: Ablation results. Qualitative ablation study across five Co3D scenes demonstrating progressive improvements from naive baseline (spatially incoherent), through pointmap conditioning (improved depth awareness), to mesh-based proximity conditioning (reduced artifacts), and finally cross-modal attention distillation (highest quality with superior consistency). Each component contributes essential capabilities that culminate in state-of-the-art performance with well-aligned modalities and enhanced realism..
B.6 Qualitative ablation
B.7 Additional scene completion results
@@
Figure 16: Qualitative results on single-view extrapolative setting. Our method can generate coherent novel view images and geometry from a single unposed reference image at extrapolative camera viewpoints, inpainting locations whose information was not given in reference images, while faithfully reconstructing the known regions.
@@
Figure 17: Qualitative comparison on two-view extrapolative setting. Our method can generate coherent novel view images and geometry from two unposed reference images at extrapolative camera viewpoints, inpainting locations not seen in reference images, while faithfully reconstructing the known regions.
@@
For two-view novel view synthesis, as demonstrated in Fig. 17, we conduct comprehensive evaluations across Co3D Reizenstein et al. (2021) datasets, demonstrating consistent performance improvements when additional reference information becomes available. The model effectively aggregates information from both reference views through our multi-view attention mechanism, resulting in more accurate geometry estimation and higher-fidelity image synthesis. Our architecture’s ability to seamlessly handle two-view inputs during inference, despite being trained on this configuration, validates the effectiveness of our correspondence conditioning and attention aggregation strategies.
B.8 Analysis on the number of input viewpoints
Our model conducts aggregated attention to generate novel views from reference images, it can receive an arbitrary number of input viewpoints for generation, as an additional reference image corresponds to simply concatenating additional reference viewpoint’s features within our attention architecture. To demonstrate this, in Figure 18, we increase the number of reference viewpoints for a model trained at 2-viewpoint setting and analyze its effects in both image quality and geometric accuracy: the results demonstrate even without being trained on the given number of inputs, our model benefits strongly from additional viewpoints, showing the generalization capability of our aggregated attention architecture to various number of input reference viewpoints.
Num. of viewpoints 1 2 3 4 5 6 7
PSNR 14.62 16.43 18.65 19.17 22.90 23.11 23.45
Table 8: Quantitative analysis regarding number of viewpoints. Our quantitative analysis conducted on Co3D dataset shows that our model’s multi-view aggregated attention enables it to generalize to an arbitrary number of reference images, with performance consistently increasing with additional reference images.
Figure 18: Analysis on number of reference viewpoints. Our model’s multi-view aggregated attention enables it to generalize to an arbitrary number of reference images, with performance consistently increasing with additional reference images.
Appendix C The use of large language models
Writing assistance. In this work, large language models (LLMs) were used exclusively for grammatical refinement and polishing of the manuscript’s writing. All usages of LLMs were only sentence-level or lower, proofreading for removal of grammatical errors, or rephrasing of certain phrases to enhance the conciseness or academic flair of the sentences originally written by the authors. All technical content, research ideas, experimental design, analysis, and scientific conclusions were developed independently by the authors without LLM assistance. The LLMs did not contribute to research ideation, methodology development, or interpretation of results. We take full responsibility for all content in this paper, including any LLM-assisted grammatical improvements.